At what angle will 510-nm light produce a second-order maximum when falling on a grating whose slits are 1.00 x 10-3 cm apart?

1
Expert's answer
2022-01-13T09:36:15-0500

Solution:

"\lambda=510\ nm=510\times10^{-9}\ m, 
\\m=2,
\\d=1\times10^{-3}\ cm=10^{-5}\ m"

We know that,

"d\sin\theta=m\lambda
\\ \therefore \theta=\sin^{-1}(\dfrac{m\lambda}{d})
\\ \Rightarrow \theta=\sin^{-1}(\dfrac{2\times510\times10^{-9}}{10^{-5}})
\\ \Rightarrow \theta=\sin^{-1}(0.102)
\\ \Rightarrow \theta=5.85\degree"